OpenStudy (snitchseeker1496):

derive (dx/dϴ) the given equation : x = Vcosϴ[(Vsinϴ-(V^(2)(sinϴ)^(2)+2gy)^(1/2))/g]

zepdrix (zepdrix):

Hey Snitch! Just so I understand your equation properly, is this how it should be formatted?\[\large\rm x=V \cos \theta\frac{Vsin \theta- \sqrt{V^2 \sin^2 \theta+2gy}{}}{g}\]

OpenStudy (snitchseeker1496):

\[x = Vcosϴ [\frac{ Vsinϴ+\sqrt{V^2(\sinϴ)^2+2gy} }{ g }]\]

OpenStudy (snitchseeker1496):

sorry it's supposed to be Vsinϴ+- in the numerator inside the bracket

zepdrix (zepdrix):

V, g constant, is y constant? or some function of x and theta or something?

OpenStudy (snitchseeker1496):

constant

OpenStudy (snitchseeker1496):

i made a mistake i'm so sorry it should be -2gy

zepdrix (zepdrix):

\[\large\rm x=V \cos \theta\left[\frac{Vsin \theta\pm\sqrt{V^2 \sin^2 \theta-2gy}}{g}\right]\] I suppose we could apply product rule, ya?\[\rm \color{royalblue}{(V \cos \theta)'}\left[\frac{Vsin \theta\pm\sqrt{V^2 \sin^2 \theta-2gy}}{g}\right]+(V \cos \theta)\color{royalblue}{\left[\frac{Vsin \theta\pm\sqrt{V^2 \sin^2 \theta-2gy}}{g}\right]'}\]

zepdrix (zepdrix):

Ahh darn, I thought that would fit on the page lol

zepdrix (zepdrix):

Hopefully you get the idea though, blue is the stuff we need to differentiate.

zepdrix (zepdrix):

The first term should be pretty straight forward. The other one will require a little work. What do you think? :d

OpenStudy (snitchseeker1496):

\[\frac{ dx }{ dϴ } = (\frac{ Vsinϴ-\sqrt{V^2\sin^2ϴ-2gy} }{ g })(-Vsinϴ) + (Vcosϴ)(\frac{ Vcosϴ-\sqrt{V^22\cosϴ\sinϴ-2gy} }{ g}\]

OpenStudy (snitchseeker1496):

shoult it be like this?

OpenStudy (snitchseeker1496):

oop it should be +/-

OpenStudy (snitchseeker1496):

The last part was cut, V^(2)2cosϴsinϴ-2gy

zepdrix (zepdrix):

\[\rm \color{orangered}{(-Vsin \theta )}\left[\frac{Vsin \theta\pm\sqrt{V^2 \sin^2 \theta-2gy}}{g}\right]+(V \cos \theta)\color{royalblue}{\left[\frac{Vsin \theta\pm\sqrt{V^2 \sin^2 \theta-2gy}}{g}\right]'}\]Your first term looks good. Let's drop it for now so it's easier to fit the other one on the page. So we have to take care of this,\[\large\rm (V \cos \theta)\color{royalblue}{\left[\frac{Vsin \theta\pm\sqrt{V^2 \sin^2 \theta-2gy}}{g}\right]'}\]

zepdrix (zepdrix):

Let's pull the 1/g to the front so it's even easier to read,\[\large\rm \left(\frac{V}{g} \cos \theta\right)\color{royalblue}{\left[Vsin \theta\pm\sqrt{V^2 \sin^2 \theta-2gy}\right]'}\]

zepdrix (zepdrix):

Your derivative of sine looks ok. Square root takes a little bit more effort though. You can convert to 1/2 power and apply power rule if you feel more comfortable doing that. But honestly, square root derivative shows up so often that it's worth remember its derivative,\[\large\rm \frac{d}{dz}\sqrt{z}=\frac{1}{2\sqrt z}\]The derivative of square root is one over two square roots.

OpenStudy (snitchseeker1496):

oh right i forgot about the square root

zepdrix (zepdrix):

\[\large\rm \left(\frac{V}{g} \cos \theta\right)\color{orangered}{\left[Vcos\theta\pm\frac{1}{2\sqrt{V^2 \sin^2 \theta-2gy}}\right]}\]So this would be our first step in dealing with the square root. Derivative of square root is one over two square roots. But we need chain rule, ya?

zepdrix (zepdrix):

Chain rule,\[\large\rm \left(\frac{V}{g} \cos \theta\right)\color{orangered}{\left[Vcos\theta\pm\frac{\color{royalblue}{\left(V^2\sin^2\theta-2gy\right)'}}{2\sqrt{V^2 \sin^2 \theta-2gy}}\right]}\]We have to multiply by the derivative of the inner function.

zepdrix (zepdrix):

I hope the color scheme is making sense. I'm using blue when we need to take a derivative, and orange when we've finished taking a derivative.

OpenStudy (snitchseeker1496):

yes thank you for helping me :)

OpenStudy (snitchseeker1496):

oh the derivative of the blue should be V^(2)2sinϴcosϴ?

OpenStudy (snitchseeker1496):

\[V^22\sinϴ\cosϴ\]?

zepdrix (zepdrix):

\[\large\rm \left(\frac{V}{g} \cos \theta\right)\color{orangered}{\left[Vcos\theta\pm\frac{\left(2V^2\sin\theta\cos\theta\right)}{2\sqrt{V^2 \sin^2 \theta-2gy}}\right]}\]Mmm yes very good. The constant on the end differentiates to 0.

zepdrix (zepdrix):

From this point, it depends how much simplifying your teacher expects you to do. It could end up being rather burdensome if you try to hard :P

OpenStudy (snitchseeker1496):

He says we have to simplify it :(

zepdrix (zepdrix):

Let's ummm..... thinking.... where to start..

OpenStudy (snitchseeker1496):

oh we have to equate dx/dϴ = 0 then we must find ϴ

zepdrix (zepdrix):

Oh oh oh, Ok that makes things much easier then.

zepdrix (zepdrix):

Lemme write down what we have so far.\[\rm -\frac{V}{g}\sin \theta\left(Vsin \theta\pm\sqrt{V^2\sin^2\theta-gy}\right)+\frac{V}{g}\cos \theta\left(V \cos \theta\pm\frac{2V^2\sin \theta \cos \theta}{2\sqrt{V^2\sin^2\theta-gy}}\right)\]Yes?

OpenStudy (snitchseeker1496):

yup

zepdrix (zepdrix):

Multiply both sides by g, divide by V,\[\rm -\sin \theta\left(Vsin \theta\pm\sqrt{V^2\sin^2\theta-gy}\right)+\cos \theta\left(V \cos \theta\pm\frac{2V^2\sin \theta \cos \theta}{2\sqrt{V^2\sin^2\theta-gy}}\right)\]

zepdrix (zepdrix):

Add the sine term to the other side,\[\rm \sin \theta\left(Vsin \theta\pm\sqrt{V^2\sin^2\theta-gy}\right)=\cos \theta\left(V \cos \theta\pm\frac{2V^2\sin \theta \cos \theta}{2\sqrt{V^2\sin^2\theta-gy}}\right)\]

zepdrix (zepdrix):

Grr this problem is a little annoying :) lol

zepdrix (zepdrix):

Oh let's get rid of the 2 in the numerator and denominator on our last term,\[\rm \sin \theta\left(Vsin \theta\pm\sqrt{V^2\sin^2\theta-gy}\right)=\cos \theta\left(V \cos \theta\pm\frac{V^2\sin \theta \cos \theta}{\sqrt{V^2\sin^2\theta-gy}}\right)\]I see a bunch of different paths we could take from here. Maybe it's just best to distribute the sine and cosine.

zepdrix (zepdrix):

\[\rm V \sin^2\theta\pm\sin \theta\sqrt{V^2\sin^2\theta-gy}=V \cos^2\theta\pm\frac{V^2\sin \theta \cos^2\theta}{\sqrt{V^2\sin^2\theta-gy}}\]

zepdrix (zepdrix):

To be honest, I'm not sure if this is going to be the best way to do this, but I'm going down a road and we'll see where it leads us :)

zepdrix (zepdrix):

Subtracting cos squared to the left side, adding the other term over to the right,\[\rm V \sin^2\theta-V \cos^2\theta=\pm\frac{V^2\sin \theta \cos^2\theta}{\sqrt{V^2\sin^2\theta-gy}}\mp\sin \theta\sqrt{V^2\sin^2\theta-gy}\]Applying our Cosine Double Angle Formula,\[\rm V \cos2\theta=\pm\frac{V^2\sin \theta \cos^2\theta}{\sqrt{V^2\sin^2\theta-gy}}\mp\sin \theta\sqrt{V^2\sin^2\theta-gy}\]

zepdrix (zepdrix):

Common denominator gives us,\[\rm V \cos2\theta=\pm\frac{V^2\sin \theta \cos^2\theta\mp\sin \theta(V^2\sin^2\theta-gy)}{\sqrt{V^2\sin^2\theta-gy}}\]

zepdrix (zepdrix):

Hmm not sure what to do from here :( What is the purpose of all of this? We're trying to find critical/stationary points for our function x?

OpenStudy (snitchseeker1496):

Ftom a engineering mechanics problem, our prof says we have to find theta in order to find Rmax in our problem. I'm just having a hard time in deriving this equation.

OpenStudy (snitchseeker1496):

Thank you for helpin me :)

zepdrix (zepdrix):

I think I need a math break, brain will esplode >.< I can't figure how to find theta from this mess. Maybe looking at it later will help. Ya hopefully that at least heads you in the proper direction D:

OpenStudy (snitchseeker1496):

Thank you :)